OpenStudy (yunomi):

Please help ^3^ 1. A group of 11 people are going to the movies one Saturday. Adult's tickets cost $12.50 each and children's tickets cost $7.75 each. If the total cost for admission to the movies is $109.00 how many children attended the movies? a.4 b.5 c.6 d.7

OpenStudy (3mar):

may I help?

OpenStudy (yunomi):

Yes please

OpenStudy (3mar):

Thank you

OpenStudy (3mar):

This kind of problem is translated easily to number of functions/equations has the same number of variables.

OpenStudy (3mar):

So what do you think of the first given?

OpenStudy (yunomi):

I don't know >.>

OpenStudy (3mar):

No problem! Let's say that: 11 people will attend some of them adult (x) and others are children (y).

OpenStudy (3mar):

so the sum of those 11is \[x+y=11............(I)\]

OpenStudy (yunomi):

okay

OpenStudy (3mar):

next given?

OpenStudy (yunomi):

?

OpenStudy (3mar):

Try with me, don't hesitate, if even get wrong. How can you translate the second given? WHat are the variables available?

OpenStudy (yunomi):

12..

OpenStudy (3mar):

12 what?

OpenStudy (3mar):

How much doe it cost for adult?

OpenStudy (3mar):

and how many adult?

OpenStudy (yunomi):

12.50 each

OpenStudy (3mar):

good: 12.50 times what?

OpenStudy (yunomi):

1? 11?

OpenStudy (yunomi):

11

OpenStudy (yunomi):

11???????????????????????

OpenStudy (3mar):

every adult will pay 12.5 and there are x adult then : 12.5x ok?

OpenStudy (3mar):

The second part is similar: number of children: y and every child pays 7.75 then: 7.75y

OpenStudy (3mar):

These two parts equal the total amount of money 109$, which is \[12.5x+7.75y=109 .............(II)\]

OpenStudy (yunomi):

okay

OpenStudy (yunomi):

then waht

OpenStudy (3mar):

then you should to solve these two simple equation together. Did you take that topic is your study before?

OpenStudy (yunomi):

solve for what how

OpenStudy (3mar):

so you did not take that lesson! what grade are you in if I may know?

OpenStudy (yunomi):

2nd

OpenStudy (3mar):

so how did you follow with me? ARE YOU IN THE 2ND GRADE?

OpenStudy (yunomi):

12th

OpenStudy (yunomi):

2. Cheryl has a taxable income of $56,742 what is her tax liability

OpenStudy (3mar):

That makes sense!

OpenStudy (3mar):

\[x+y=11\] \[12.5x+7.75y=109\] solve : x=5 y=6 adults are five and 6 children.

OpenStudy (yunomi):

is the answer c?

OpenStudy (yunomi):

c?

OpenStudy (3mar):

Correct! C

OpenStudy (yunomi):

really? yay

OpenStudy (3mar):

Yes

OpenStudy (yunomi):

Can you help me with another

OpenStudy (3mar):

Before go to another post, did you get the idea?

OpenStudy (yunomi):

I think so

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (yunomi):

Help with another? please?

OpenStudy (yunomi):

1. Gabriel needs to purchase baseballs and bats for his youth baseball league. He'll need at least 20 peices of equaipment. but he only has $525 to spend. Baseballs cost $20 each and bats cost $35 each. If Gabriel wants to purchase the largest amount of equipment and still stay within his budget , which combination of baseballs and bats is optimal?

OpenStudy (yunomi):

a. (19,1) b.(21,3) c.(14,7) d.(16,4)

OpenStudy (yunomi):

help please

OpenStudy (3mar):

Sorry! It was an emergency call.